FEL Upgrade Project Weekly Brief

February 11-15, 2002

 

 

Highlights:

Significant progress on the gun occurred with the chamber going out for vacuum firing and the support tube getting the inner ring welded.

The first production coil for the optical chicane dipoles was successfully produced.

Management:

F. Dylla visited the Wellman Lab of Photomedicine on Feb. 13 to present a seminar and discuss potential collaborations on FEL applications to bio-medicine. He visited the Engineering School at Northeastern University on Feb.15 to present a seminar on materials processing with the FEL.

For the high power injector work, proposed 750 MHz injector cavity shapes were received from AES for review and analysis for multipacting.

Discussions were held on approaches for wiggler replacement for 1 micron output optimization. UV cavity approaches were the subject of a separate meeting. Analysis and discussions in both of those areas continues.

Efforts are underway to incorporate each WBS element's installation schedule as based on expected delivery and checkout of hardware into a master installation plan. Also to be incorporated are outside activities that relate to resource loading such as CEBAF operation schedule, etc.

A DOE Occurrence Report was produced to close out reporting requirements on an electrical shock incident two weeks ago.

WBS 3 (Beam Physics):

Work continued on machine modeling, commissioning plan, and evaluation of magnet field quality (both simulated [dipoles] and measured [quadrupoles])

WBS 4 (Injector):

The gun chamber is out to VacMet for vacuum firing. The extractor cart drawings were signed off and submitted for bids. The support tube had the inner ring welded in and is having the oxide regions polished out. Part fit up and fine tuning continues.

Gun HVPS - Waiting for the new tank. The design of the mechanism for the new combined Conditioning & Running resistors is still in progress. The resistive inks and protective coatings have been tested as several sample resistors were made.

Gun HVPS Tank - The contract for the tank has not yet been placed.

PARMELA simulations of the transverse and longitudinal emittance evolution have been performed for 135 pC and 500 kV. So far the simulation has been run from the gun up to IMS0F03. The effect of the number of particles used for the simulations is under study. Similar simulations of the bunch length (in ps and mm) have been performed from the GUN up to MDU0F03 to provide information for ultra-fast laser imaging of the electron bunch. Efforts to implement PARMELA on PC continue.

Previously taken data of implanted electrode has been fitted with Field Emission equation. Results of the simulation show that the field enhancement factor beta ~ 260, the emitter height ~ 300 nm, emitter tip radius ~ 1 nm and the work function (which is unknown for the implanted coating) is very similar to that of Stainless Steel ~4.25 eV.

A new wafer is under vacuum waiting to be hydrogen cleaned in preparation for field emission characterization of its surface.

WBS 6 (RF):

Quarter Klystrons - The klystron cart drawings are being modified for the water manifold changes. They are due next week. Two additional carts will then be built.

Quarter HVPS - Wiring of the HVPS continues. Drawings are being kept up to date with any changes or corrections. The Electricians have started wiring the 480 VAC sections.

WBS 8 (Instrumentation):

The majority of the extra cabling in the vault has been removed. We are looking at different methods of cable routing for the new system to make better use of the space available and ease tray loading. Cable tray modifications are complete for the time being more trays will be added once component location has been determined.

A test fixture for the new CCD cameras has been built so each camera can be tested as it is built and will be ready for installation. A written procedure has been developed that outlines both camera assembly and testing.

Parts are on order for the new 8-channel OMS stepper motor chassis with fabrication expected to begin next month.

Initial mapping of the magnet power supplies is underway. Trim channels are also being designated as required. Cable routing for these items will be starting soon.

The MPS VME interface card check print is complete and design work continues in switching the Beam Loss Monitors from CAMAC to VME based.

 

WBS 9 (Transport):

Dipoles

At Magnet Enterprises International in Oakland CA. the first production coil was potted successfully. The second coil of the first magnet is wound and is in preparation for potting.

o Return legs are being annealed. Top and bottom slabs are still rough machining. Other parts are on order or are being worked on.

Injector Dipoles (DU/DV)

o Contracts were let to Wang NMR for coils and Assembly and to Master Machine for Cores and Parts for the Injector Dipoles.

o DULY Research completed their checking the quality of the field of the Small Injector Dipole (GV) and has a few changes that will have to be incorporated in the core parts in a Revision A.

Arc Dipoles (GY, GX, GQ)

o Electronic procurement of Bend (GX) and Reverse Bend (GQ) Dipoles continues.

o Advanced Energy Systems (AES) completed their analysis of tall three magnets at low excitation for low energy for UV as well as IR. Results indicate we will have to shorten the dipoles by up to 5 mm for the two lower energy nominal cases. For high energy, we will have to add back some shim pieces to the face of the dipoles.

o Wang NMR of CA is continuing to insulate the conductor for the 180 Dipole Coils and design of the winding form and potting mold. They have started building the winding form for the path Length Correctors (GG Coils) that fit into the faces of these magnets.

o Bosma Machine has started rough machining of the core pieces.

Quadrupoles

3 inch quad (QX) Magnet Measurement Refinement.

o We are concentrating on getting the magnet measurement stand qualified to measure the quadrupoles. We are confronted with data that has scatter of the same order as that of the tolerance in the specification for field quality. We are assessing how to measure the magnets and handle the data to reduce the scatter.

o Additional mechanical measurements show the machined contact surfaces are tied to the parting plane notches within at least half of the one mradian roll specification. (And as reported last week, those parting plane surfaces are within .001 inch of the pole tip edges.) This indicates that we will be able to reliably use these surfaces for accurate alignment.

o Additional magnetic measurements show that by offsetting the centerline of the quadrupole by incremental distances on the test stand there are corresponding changes to the dipole term. When this term is cranked through our algorithm that calculates the offset, the results are over double the real offset and are not along the correct plane. However, each incremental calculated position is along a straight line radiating from an axis. Over the next week, we plan to refine the algorithm so that the real and calculated offsets correspond. The goal is to use the dipole term characterize the center of the quadrupole too well within the alignment specifications.

Trim Quad (QT)

o Milhous Control of Virginia completed trial windings and will be ready to start actual winding next week.

Sextupole (SF)

o DULY has a magnetic model that is within specification but has poles that are less wide that is ideal for fabrication. They are going through some iteration to thicken the pole.

Octupole (OT)

o Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

o AES continues drawing up the chambers through out the arcs. Several mismatches in the data we supplied them for chamber position resulted in a need to reestablish the baseline.

o We are finishing of the regions before the first arc and after the second arc.

WBS 10 (Wiggler):

I finished analyzing the commissioning data for the second wiggler. The corrector field seems to be a bit stronger than it should be, leading to a 3 G DC component in the field. The initial commissioning values of the corrector fields will be corrected to account for this. We are now taking data on the vertical field to put a lower limit on the DC remanent in that direction. Dispersion section commissioning data collection will commence next week.

WBS 11 (Optics):

We continued our testing of the 3" HR deformable mirror assembly (DMA) in the mirror test stand (MTS). We believe the laser in the laser interferometer is beginning to fail, and expect a service call next week.

The revised drawings for the optical cavity assemblies were released yesterday for a final check.caught a few more changes that are being incorporated, this will delay release of the bid package by one week. After receiving budgetary estimates from the leadscrew vendor, we have modified the drawings; they are being redone and e should be able to go out for procurement next week. We are also pressing forward with plans to finish specifications for the outcouplers and get them procured. Design of the insertable mirror assemblies continues. We have continued work on the OCMMS, and have begun looking into low-cost ways to get the OCMMS HeNe beam size into EPICS. Fortunately, it appears we can use solutions developed outside the lab to do this.

We have submitted a requisition for the ultrafast laser that will be used to set the cavity length and perform other FEL diagnostics.

Other Activities

We have received a budgetary estimate for a new drive laser oscillator and timing electronics. The price looks reasonable. The current drive laser was started and checked out, it is operating well.

 

 

 

 

--George/Fred/dmag